are the two equations -6+y=2x and 2y-4x=12 dependant?
Answer by ewatrrr(24785)

 
Hi,
are the two equations -6+y=2x 0r y -2x = 6 and 2y-4x=12 dependant?
y -2x = 6 and
2y -4x = 12
Multiplying first by -2 and adding to 2nd EQ
-2y +4x = 12 and
2y -4x = 12
0 = 0
A. Yes, because they have the same graph y = 2x+6
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions. ****